# Denoising Task Difficulty-based Curriculum for Training Diffusion Models

### v2

@@ 1 Introduction
Diffusion-based generative models [18, 52, 57] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image [8], video [17, 15], and 3D [62, 35] generation. Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 40], class-conditional [8], and text-conditional scenarios [2, 47], demonstrate their significant impact. Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
@@
Diffusion models [18, 57] are designed to reverse the corruption of the data through the learning process at different noise levels and over multiple timesteps. Recent works have delved into the learning of diffusion models across various noise levels and timesteps, revealing different stages of diffusion models. For example, Choi et al. [5] observe that when a diffusion model performs a denoising task from large to small timestep, it first generates coarse features, then gradually generates perceptually rich content, and later refines the details. Similar observation is also identified in text-to-image diffusion models [2]. Besides this aspect, various studies have further explored the learning of diffusion models across timesteps and noise levels, elucidating their transition from denoising to generative functionalities [6], modular attributes [67], frequency characteristics [66, 31], trajectories [41], affinity [11], and variations of targets [64].
@@
These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models. Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 64], architectures [31, 2], accelerated sampling [41], representations [67], and guidance [12]. Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering insights and achieving unprecedented advancements in their capabilities.
@@
In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising. Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse. Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult [18, 14], while others argue that smaller timesteps pose greater difficulties [22, 9, 25]. The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
@@
Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 27, 4, 61, 45], into the training process of diffusion models. Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with ascending levels of difficulty, moving from clusters categorized by higher to lower timesteps. After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 57, 19] to reach the convergence point. Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
Building upon this foundation, our curricular approach offers several notable advantages: 1) Improved Performance and 2) Faster Convergence: By leveraging the inherent benefits of curriculum learning, our method significantly enhances the quality of generation and the speed of convergence. 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models. Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14, 11, 23].
Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks. These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [24], ImageNet [7], and MS-COCO [33]. By integrating our curriculum learning strategy into architectures— DiT [44], EDM [22], and SiT [38]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
@@ 2.1 Diffusion Models
Diffusion models [18, 52, 57] are a group of generative models that create samples by utilizing a learned denoising process to noise. Several works have focused on improving diffusion models in various aspects, including model architectures [43, 8, 42], sampling speed [53, 37, 34], training objectives [14, 5, 11, 26, 38]. These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models. P2 [5] under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps. Ediff-I [2] observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions. They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process. Moreover, various works have also investigated these aspects related to timesteps [6, 67, 66, 31, 41, 11, 64] (detailed illustrations can be found in Appendix A). While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty. Furthermore, we leverage these observations to propose a curriculum learning approach.
@@ 2.2 Denoising Difficulties on Diffusion Models
Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored. Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
Certain studies consider denoising at larger noise levels and timesteps to be more difficult [18, 14], the focus is on the challenges associated with reconstructing data from substantial noise. For instance, Hang et al. [14] articulate that while smaller timesteps (approaching zero) may require straightforward reconstructions, such strategies become less effective at higher noise levels or in larger timesteps. Similarly, Ho et al. [18] elucidate that their approach de-emphasizes loss terms at smaller timesteps to prioritize learning on the more challenging tasks at larger timesteps, thereby enhancing sample quality. Conversely, other studies argue that earlier timesteps or lower noise levels also present significant challenges. Karras et al. [22] suggest that detecting noise at low levels is challenging due to its minimal presence. Also, Kim et al. [25] illustrate the increasing difficulty and high variance in score estimation as timesteps approach zero, disturbing stable training of models. In line with these observations, Dockhorn et al. [9] build upon insights of [25], acknowledging the complexities at near zero timesteps, where score becomes highly complex and potentially unbounded.
@@ 2.3 Curriculum Learning
Curriculum learning [3, 13, 27], inspired by human learning patterns, is a method of training models in a structured order, starting with easier tasks [45] or examples [3] and gradually increasing difficulty. As pointed out by [3], curriculum learning formulation can be viewed as a continuation method [1], which starts from a smoother objective and gradually transformed into a less smooth version until it reaches the original objective function. Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly [13, 27, 4, 61].
Curriculum learning primarily comprises two components: a curriculum scoring function, measuring the difficulty of tasks or examples, and a pacing function, modulating the speed of the curriculum progress. Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [58]. Recently, various works employ models to measure difficulty, including confidence of pre-trained models [13] and the loss of the current models [27]. For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 63]. There are various forms of this and they can be generally represented as a function of training iteration [13, 63]. Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training [29, 21].
In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [59, 65]. Tang et al. [59] sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets. Similarly, Yang et al. [65] order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets. In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps. Also, while consistency models [55, 54] adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
@@ 3 Preliminaries
In this section, we provide the necessary background on diffusion models [18, 52, 57]. Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$. The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise. This can be formulated as a stochastic differential equation (SDE) [57] as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
@@
where $\bar{\bm{w}}_{t}$ and $\text{d}\bar{t}$ denote the reverse-time Wiener process and the infinitesimal reverse-time, respectively, with the actual data score $\nabla\log p_{t}(\bm{x}_{t})$. In most cases, a neural network $\epsilon_{\theta}$ having parameter $\theta$ is utilized to approximate this score function by learning the denoising tasks for each timestep $t$ from score matching loss $\mathcal{L}$ [56]:
@@
where $\omega(t)$ is loss weights for $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial timestep 00 to $t$. This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$. This is regularly denoted as $\epsilon$-prediction parameterization [18, 20], and several other parameterizations including $F$-prediction [22, 26], score-prediction [57] and velocity-prediction [38] have been proposed.
@@ 4 Observations
4.1 Analysis on the Task Difficulty in terms of Convergence
In this study, we analyze the convergence speed of loss and denoising performance across timesteps. To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT [20] for $\epsilon$-prediction, EDM [22] for $F$-prediction, SiT [38] for velocity prediction. Detailed descriptions of the experimental setups are provided in Appendix B.
@@ 4.1 Analysis on the Task Difficulty in terms of Convergence
We also delve deeper into a convergence of denoising performance according to timesteps with 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$. To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in [11]. Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on denoising tasks throughout the training process. Figures 1(d)-1(f) depict this convergence. They illustrate that, similar to loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed across the DiT, EDM, and SiT. These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.

### v3

@@ 1 Introduction
Diffusion-based generative models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021) have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image (Dhariwal & Nichol, 2021), video (Ho et al., 2022a, Harvey et al., 2022), and 3D (Woo et al., 2023, Liu et al., 2023b) generation. Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional (Karras et al., 2022, Nichol & Dhariwal, 2021), class-conditional (Dhariwal & Nichol, 2021), and text-conditional scenarios (Balaji et al., 2022, Ramesh et al., 2022), demonstrate their significant impact. Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
@@
Diffusion models (Ho et al., 2020, Song et al., 2021) are designed to reverse the corruption of the data through the learning process at different noise levels and over multiple timesteps. Recent works have delved into the learning of diffusion models across various noise levels and timesteps, revealing different stages of diffusion models. For example, Choi et al. (Choi et al., 2022) observe that when a diffusion model performs a denoising task from large to small timestep, it first generates coarse features, then gradually generates perceptually rich content, and later refines the details. Similar observation is also identified in text-to-image diffusion models (Balaji et al., 2022). Besides this aspect, various studies have further explored the learning of diffusion models across timesteps and noise levels, elucidating their transition from denoising to generative functionalities (Deja et al., 2022), modular attributes (Yue et al., 2024), frequency characteristics (Yang et al., 2023b, Lee et al., 2023), trajectories (Pan et al., 2024), affinity (Go et al., 2023a), and variations of targets (Xu et al., 2023).
@@
These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models. Specifically, these insights are incorporated into their method design in various works, including loss functions (Hang et al., 2023, Xu et al., 2023), architectures (Lee et al., 2023, Balaji et al., 2022), accelerated sampling (Pan et al., 2024), representations (Yue et al., 2024), and guidance (Go et al., 2023b). Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering insights and achieving unprecedented advancements in their capabilities.
@@
In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising. Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties (Li et al., 2023, Balaji et al., 2022), yet a detailed exploration of these variances remains sparse. Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult (Ho et al., 2020, Hang et al., 2023), while others argue that smaller timesteps pose greater difficulties (Karras et al., 2022, Dockhorn et al., 2021, Kim et al., 2022). The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
@@
Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature (Hacohen & Weinshall, 2019, Kong et al., 2021, Chang et al., 2021, Wang et al., 2020, Pentina et al., 2015), into the training process of diffusion models. Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with ascending levels of difficulty, moving from clusters categorized by higher to lower timesteps. After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training (Ho et al., 2020, Song et al., 2021, Ho & Salimans, 2022) to reach the convergence point. Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
Building upon this foundation, our curricular approach offers several notable advantages: 1) Improved Performance and 2) Faster Convergence: By leveraging the inherent benefits of curriculum learning, our method significantly enhances the quality of generation and the speed of convergence. 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models. Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting (Choi et al., 2022, Hang et al., 2023, Go et al., 2023a, Karras et al., 2023).
Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks. These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ (Karras et al., 2019), ImageNet (Deng et al., 2009), and MS-COCO (Lin et al., 2014). By integrating our curriculum learning strategy into architectures— DiT (Peebles & Xie, 2022), EDM (Karras et al., 2022), and SiT (Ma et al., 2024)—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021) are a group of generative models that create samples by utilizing a learned denoising process to noise. Several works have focused on improving diffusion models in various aspects, including model architectures (Park et al., 2024b, Dhariwal & Nichol, 2021, Park et al., 2024a), sampling speed (Song et al., 2020, Lu et al., 2022, Liu et al., 2023a), training objectives (Hang et al., 2023, Choi et al., 2022, Go et al., 2023a, Kingma & Gao, 2023, Ma et al., 2024). These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models. P2 (Choi et al., 2022) under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps. Ediff-I (Balaji et al., 2022) observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions. They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process. Moreover, various works have also investigated these aspects related to timesteps (Deja et al., 2022, Yue et al., 2024, Yang et al., 2023b, Lee et al., 2023, Pan et al., 2024, Go et al., 2023a, Xu et al., 2023) (detailed illustrations can be found in Appendix A). While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty. Furthermore, we leverage these observations to propose a curriculum learning approach.
@@ 2.2 Denoising Difficulties on Diffusion Models
Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored. Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties (Li et al., 2023, Balaji et al., 2022), and there have been conflicts regarding these difficulties between previous works.
Certain studies consider denoising at larger noise levels and timesteps to be more difficult (Ho et al., 2020, Hang et al., 2023), the focus is on the challenges associated with reconstructing data from substantial noise. For instance, Hang et al. (Hang et al., 2023) articulate that while smaller timesteps (approaching zero) may require straightforward reconstructions, such strategies become less effective at higher noise levels or in larger timesteps. Similarly, Ho et al. (Ho et al., 2020) elucidate that their approach de-emphasizes loss terms at smaller timesteps to prioritize learning on the more challenging tasks at larger timesteps, thereby enhancing sample quality. Conversely, other studies argue that earlier timesteps or lower noise levels also present significant challenges. Karras et al. Karras et al. (2022) suggest that detecting noise at low levels is challenging due to its minimal presence. Also, Kim et al. (Kim et al., 2022) illustrate the increasing difficulty and high variance in score estimation as timesteps approach zero, disturbing stable training of models. In line with these observations, Dockhorn et al. (Dockhorn et al., 2021) build upon insights of (Kim et al., 2022), acknowledging the complexities at near zero timesteps, where score becomes highly complex and potentially unbounded.
@@ 2.3 Curriculum Learning
Curriculum learning (Bengio et al., 2009, Hacohen & Weinshall, 2019, Kong et al., 2021), inspired by human learning patterns, is a method of training models in a structured order, starting with easier tasks (Pentina et al., 2015) or examples (Bengio et al., 2009) and gradually increasing difficulty. As pointed out by (Bengio et al., 2009), curriculum learning formulation can be viewed as a continuation method (Allgower & Georg, 2003), which starts from a smoother objective and gradually transformed into a less smooth version until it reaches the original objective function. Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly (Hacohen & Weinshall, 2019, Kong et al., 2021, Chang et al., 2021, Wang et al., 2020).
Curriculum learning primarily comprises two components: a curriculum scoring function, measuring the difficulty of tasks or examples, and a pacing function, modulating the speed of the curriculum progress. Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images (Bengio et al., 2009) or the length of sequences (Spitkovsky et al., 2010). Recently, various works employ models to measure difficulty, including confidence of pre-trained models (Hacohen & Weinshall, 2019) and the loss of the current models (Kong et al., 2021). For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression (Hacohen & Weinshall, 2019, Wu et al., 2020). There are various forms of this and they can be generally represented as a function of training iteration (Hacohen & Weinshall, 2019, Wu et al., 2020). Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training (Kumar et al., 2010, Jiang et al., 2014).
In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks (Tang et al., 2023, Yang et al., 2023a). Tang et al. (Tang et al., 2023) sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets. Similarly, Yang et al. (Yang et al., 2023a) order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets. In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps. Also, while consistency models (Song et al., 2023, Song & Dhariwal, 2023) adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
@@ 3 Preliminaries
In this section, we provide the necessary background on diffusion models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021). Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$. The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise. This can be formulated as a stochastic differential equation (SDE) (Song et al., 2021) as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
@@
where $\bar{\bm{w}}_{t}$ and $\text{d}\bar{t}$ denote the reverse-time Wiener process and the infinitesimal reverse-time, respectively, with the actual data score $\nabla\log p_{t}(\bm{x}_{t})$. In most cases, a neural network $\epsilon_{\theta}$ having parameter $\theta$ is utilized to approximate this score function by learning the denoising tasks for each timestep $t$ from score matching loss $\mathcal{L}$ (Song & Ermon, 2019):
@@
where $\omega(t)$ is loss weights for $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial timestep 00 to $t$. This object can be interpreted as a noise-matching loss in DDPM (Ho et al., 2020), which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$. This is regularly denoted as $\epsilon$-prediction parameterization (Ho et al., 2020, Jabri et al., 2022), and several other parameterizations including $F$-prediction (Karras et al., 2022, Kingma & Gao, 2023), score-prediction (Song et al., 2021) and velocity-prediction (Ma et al., 2024) have been proposed.
@@ 4 Observations
4.1 Analysis on the Task Difficulty in terms of Convergence Speed
In this study, we analyze the convergence speed of loss and denoising performance across timesteps. To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT (Jabri et al., 2022) for $\epsilon$-prediction, EDM (Karras et al., 2022) for $F$-prediction, SiT (Ma et al., 2024) for velocity prediction. Detailed descriptions of the experimental setups are provided in Appendix B.
@@ 4.1 Analysis on the Task Difficulty in terms of Convergence Speed
We also delve deeper into a convergence of denoising performance according to timesteps with 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$. To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in (Go et al., 2023a). Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score (Heusel et al., 2017), enabling us to observe the performance convergence of each model on denoising tasks throughout the training process. Figures 1(d)-1(f) depict this convergence. They illustrate that, similar to loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed across the DiT, EDM, and SiT. These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.